# Long-Tailed 3D Detection via Multi-Modal Late-Fusion

### v4

@@ 5.2.3 Analysis of Misclassification
We evaluate CenterPoint [12] and CMT [10] using the hierarchical mAP metric (mAPH), which considers semantic relationships across classes to award partial credit. In safety-critical applications, detecting but misclassifying objects (as a semantically related category) is more desirable than a missed detection (e.g., detecting but misclassifying a child as an adult is preferable to not detecting this child). Table V lists the results of applying our developed techniques including the group-free detector head, the hierarchy loss, and MMLF using the 2D detector DINO. Results reveal that (1) classes are most often misclassified as their LCA=1 siblings within coarse-grained superclasses, (2) our developed techniques significantly improve mAPH (especially on rare classes), reducing misclassification between sibling classes.
@@ 5.2.4 Analysis of Failure Cases
We visualize common failure cases of MMLF and compare them with the failure cases of TransFusion [34], an end-to-end trained multi-modal detector. We find that our MMLF fails in cases of occlusions (where there is no 3D information) and in cases where the 2D RGB detector misclassifies objects. See Fig. 10 for detailed analysis.
Figure 11: Training 2D detectors with more data (e.g., training with nuScenes + nuImages vs. nuScenes only) and using better 2D detectors (e.g., DINO vs. YOLOV7) improve performance on the proxy task of 2D detection and the downstream 3D detection achieved by our multi-modal late-fusion algorithm.
5.3 Limitations and Future Work
Our work does not directly study how addressing LT3D affects downstream perception tasks. Future work should address this limitation. Moreover, as shown in Fig. 11, simply training better 2D RGB detectors with more data provides a natural pathway for improving LT3D performance. We find that 2D detection accuracy is a strong proxy for final 3D LT3D performance. Recent works in large-scale vision language models [77, 78, 79, 80, 81] show promising zero-shot results in detecting rare classes. Identifying ways of incorporating foundation models into our late-fusion framework might improve LT3D further. In addition, our proposed late-fusion approach is unlikely to work well for objects that are visible in LiDAR but occluded in RGB images (or visible in RGB images but occluded in LiDAR, but this is unlikely because LiDARs are typically mounted higher than RGB cameras). Notably, due to LiDAR and RGB sensors not being co-located on the autonomous vehicle, some objects may be visible in one modality but occluded in another. If an object is visible in LiDAR but not in RGB, we would down-weight the confidence scores of LiDAR-based detections that not supported by RGB-based detections. In contrast, if an object is visible in RGB but not in LiDAR, our 2D late-fusion pipeline would not be able to detect this object in 3D.
@@ 6 Conclusion
We explore the problem of long-tailed 3D detection (LT3D), detecting objects not only from common classes but also from many rare classes. This problem is motivated by the operational safety of autonomous vehicles (AVs) but has broad robotic applications, e.g., elder-assistive robots [82] that fetch diverse items [83] should address LT3D. To study LT3D, we establish rigorous evaluation protocols that allow for partial credit to better diagnose 3D detectors. We propose several algorithmic innovations to improve LT3D, including a group-free detector head, a hierarchical loss that promotes feature sharing across long-tailed classes, and present a detailed exploration of multi-modal late-fusion for LT3D. We find that 2D RGB detectors are better suited for late-fusion, matching projected 3D LiDAR detections on the 2D image-plane outperforms matching 2D RGB detections inflated to 3D, and score calibration and probabilistic fusion notably improve performance. Our simple multi-modal late-fusion achieves state-of-the-art LT3D performance on the nuScenes benchmark, improving over prior works by 5.9 mAP over all long-tailed distributed classes!
Acknowledgments
This work was supported by the University of Macau (SRG2023-00044-FST), NSFC (No.62206256), CMU Center for Autonomous Vehicle Research, and NSF GRFP (DGE2140739).
A Implementation Details
We use the open-source implementations from [9] for training FCOS3D, CenterPoint, and TransFusion, and the first-party implementations for other detectors, following the training schedule proposed by each respective paper. By default, we train the 2D RGB detector with 2D bounding boxes derived from nuScenes’ 3D annotations and additionally train with 2D bounding boxes from nuImages where denoted. Our 2D RGB detectors, YOLOV7 and DINO, are pre-trained on the ImageNet [84] and COCO [55] datasets. We follow the default training recipes for both YOLOV7 and DINO respectively. We describe important implementation details below.
@@ A Implementation Details
  • Training Losses. We use the sigmoid focal loss (for recognition) [60] and L1 regression loss (for localization) below. Existing works also use the same losses but only with fine labels; we apply the loss to both coarse and fine labels. Concretely, our loss function for CenterPoint is as follows: $L=L_{HM}+\lambda L_{REG}$, where $L_{HM}=\sum_{i=0}^{C}SigmoidFocalLoss(X_{i},Y_{i})$ and $L_{REG}=|X_{BOX}-Y_{BOX}|$, where $X_{i}$ and $Y_{i}$ are the $i^{th}$ class’ predicted and ground-truth heat maps, while $X_{BOX}$ and $Y_{BOX}$ are the predicted and ground-truth box attributes. Without our hierarchical loss, $C$=18. With our hierarchical loss, $C$=22 (18 fine grained + 3 coarse + 1 object class). $\lambda$ is set to 0.25. Modifications for other detectors similarly follow.
  • Optimization. We train all LiDAR-only detectors for 20 epoch using an AdamW optimizer and a cyclic learning rate. We adopt a basic set of data augmentations, including global 3D tranformations, flip in BEV, and point shuffling during training. We train our model with 8 RTX 3090 GPUs and a batch size of 1 per GPU. The training noise (from random seed and system scheduling) is $<$ 1% of the accuracy (standard deviation normalized by the mean).
@@
B More Visualizations
C Analysis on Transformer 3D Detectors
In the main paper, we evaluate all late-fusion methods by filtering LiDAR detections from CenterPoint with different RGB detectors for fair comparison to [9]. However, our late-fusion approach generalizes to other 3D LiDAR detectors as well. To this end, we evaluate our best-performing late-fusion configuration from the main paper with LiDAR detections from BEVFusion-L [35] and CMT-L [10] in Table VI.
@@ C Analysis on Transformer 3D Detectors
TABLE VI: Transformer-based 3D LiDAR detector. We evaluate the late-fusion performance of BEVFusion-L and CMT-L (LiDAR-only) with DINO, and find that our late fusion strategy improves over BEVFusion-L by 8.2 mAP overall and CMT-L by 15.9 mAP. This suggests that our late-fusion approach generalizes across different 3D detector architectures.
@@
TABLE VII: Comparison of inference time. Our multi-modal late-fusion approach runs faster than existing methods including end-to-end multimodal detectors such as DeepInteraction. Note that although our approach runs two uni-modal detectors, we run them in parallel so that the detection stage maintains the same efficiency as CenterPoint (CP). Moreover, our late-fusion strategy introduces negligible computation in inference. We copy 3D detection performance (in mAP) from the main paper for reference. Results demonstrate that our simple multi-modal late-fusion approach resoundingly outperforms sophisticated state-of-the-art methods (e.g., DeepInteraction and TransFusion) with significantly less inference time!
FCOS3D BEVFormer CP TransFusion DeepInteraction MMLF(CP, DINO)
@@
D Inference Runtime
Table VII compares the inference runtime of our method against prior works on a single GPU. Although our multi-modal late-fusion (MMLF) methods such as MMLF(CenterPoint, DINO) run two uni-modal detectors, we run them in parallel. Therefore, our final detector maintains the same runtime as a typical uni-modal detector and is often faster than more sophisticated multimodal detectors like BEVFormer [23], TransFusion [34] and DeepInteraction [59]. However, running two detectors in parallel will require more memory usage. Importantly, late-fusion methods including CenterPoint + DINO and [9] have negligible overhead when fusing uni-modal detections, hence their runtime is the same as CenterPoint, which serves as the speed bottleneck. This work uses research-level code and can be further optimized for deployment on autonomous vehicles.
TABLE VIII: Our proposed group-free detector head architecture consistently outperforms grouping-based approaches on both the standard and LT3D  benchmarks. We note that sub-optimal grouping strategies (such as those adopted for LT3D) may yield significantly diminished performance, whereas optimized grouping strategies (such as those adopted for the standard setup) have comparable performance to the group-free approach. Note, TC is traffic-cone, CV is construction vehicle, MC is motorcycle, PP is pushable-pullable, CW is construction-worker, and PO is police-officer. We highlight classes with Medium and Few examples per class in blue.
@@ D Inference Runtime
Adult PP CW Debris Child Stroller PO EV PM All
@@
E Analysis on Class Grouping
@@
We use the class groups proposed by prior works [18, 12] for the standard benchmark and adapt this grouping for LT3D. Our proposed group-free detector head architecture consistently outperforms grouping-based approaches on both the standard and LT3D  benchmarks. We note that sub-optimal grouping strategies (such as those adopted for LT3D) may yield significantly diminished performance, whereas optimized grouping strategies (such as those adopted for the standard setup) have comparable performance to the group-free approach. The group-free approach simplifies architecture design, while also providing competitive performance.
@@
TABLE IX: Different variants achieve similar performance. We note that other methods do improve accuracy in the tail by sacrificing performance in the head, suggesting that hybrid approaches that apply different techniques for head-vs-tail classes may further improve accuracy. Unlike [87, 88] which requires a strict label hierarchy, our approach is not limited to a hierarchy.
@@
F Analysis of Hierarchical Training
@@
Classic methods train a hierarchical softmax (in contrast to our simple approach of sigmoid focal loss with both fine and coarse classes), where one multiplies the class probabilities of the hierarchical predictions during training and inference [88]. We implemented such an approach, but found the training did not converge. Interestingly, [88] shows such a hierarchical softmax loss has little impact on long-tailed object detection (in 2D images), which is one reason they have not been historically adopted. Instead, we found better results using the method from [87] (a winning 2D object detection system on the LVIS [52] benchmark) which multiples class probabilities of predictions (e.g. $P_{CAR}=P_{OBJ}*P_{CAR}$) at test-time, even when such predictions are not trained with a hierarchical softmax. We tested three variants and compared it to our approach (which recall, uses only fine-grained class probabilities at inference). Table IX compares their performance for LT3D.
TABLE X: Per-class performance on nuScenes. Our late-fusion approach achieves the highest per-class AP on 5 out of 10 classes. Compared to DeepInteraction, our approach improves construction worker by 25.2 AP, stroller by 6.8 AP, and pushable-pullable by 27.3 AP. Note, CV is construction vehicle, MC is motorcycle, PP is pushable-pullable, CW is construction-worker, and Stro. is stroller. We highlight classes with Medium and Few examples per class in blue.
Method Car Adult Truck CV Bicycle MC Child CW Stroller PP
TransFusion [34] 84.4 84.2 58.4 24.5 46.7 60.8 3.1 21.6 13.3 25.3
BEVFusion [35] 90.2 67.2 65.5 35.2 58.8 77.0 4.4 39.0 29.6 34.1
DeepInteraction [59] 84.9 85.9 63.2 35.3 64.3 76.2 6.0 30.7 30.9 30.8
CMT [10] 88.6 87.7 65.2 36.9 66.7 76.3 4.7 34.4 9.4 34.1
MMLF(CP, DINO) 86.3 86.2 60.6 35.3 70.1 75.9 8.8 55.9 37.7 58.1
TABLE XI: Evaluating 2D detection performance. We find that simply projecting 3D detections from RGB-only 3D detectors to 2D image plane yields considerably lower 2D detection mAP across all classes. However, 2D RGB detectors achieve higher performance. Intuitively, 2D RGB detectors achieve better classification performance and predict tighter fitting bounding boxes than those derived by projecting 3D detections to the image plane. This explains why fusing detections from 2D RGB detectors yields better LT3D performance than detections from 3D RGB detectors. All detectors in this table are only trained on the nuScenes train-set.

### v5

@@ 6.2.3 Analysis of Misclassification
We evaluate CenterPoint [82] and CMT [80] using the hierarchical mAP metric (mAPH), which considers semantic relationships across classes to award partial credit. In safety-critical applications, detecting but misclassifying objects (as a semantically related category) is more desirable than a missed detection (e.g., detecting but misclassifying a child as an adult is preferable to not detecting this child). Table 5 highlights the results of applying our proposed techniques including the group-free detector head, the hierarchy loss, and MMLF using the 2D detector DINO. Results reveal that (1) classes are most often misclassified as their LCA=1 siblings within coarse-grained superclasses, (2) our developed techniques significantly improve mAPH (especially on rare classes), reducing misclassification between sibling classes.
@@ 6.2.4 Analysis of Failure Cases
In Fig. 10, we visualize common failure cases of MMLF and compare them with the failure cases of TransFusion [2], an end-to-end trained multi-modal detector. Both TransFusion and our MMLF fail in cases of occlusions (where there is no 3D information) and in cases where the 2D RGB detector misclassifies objects.
Figure 11: Training 2D detectors with more data (e.g., training with nuScenes + nuImages vs. nuScenes only) and using better 2D detectors (e.g., DINO vs. YOLOV7) improve performance on the proxy task of 2D detection and the downstream 3D detection achieved by our multi-modal late-fusion algorithm. See Appendix I for further analysis on the impact of 2D detection quality.
6.2.5 Analysis of Using Strong 2D Detectors
As shown in Fig. 11, simply training better 2D RGB detectors with more data provides a natural pathway for improving LT3D performance. It suggests a strong correlation between 2D detection accuracy and the final 3D LT3D performance. This is an exciting signal, because, as more and more powerful foundational 2D detectors become publicly available [45, 60, 56] , our MMLF framework can utilize these more advanced 2D detectors to further improve LT3D performance.
@@ 7 Conclusion
We explore the problem of long-tailed 3D detection (LT3D), detecting objects not only from common classes but also from many rare classes. This problem is motivated by the operational safety of autonomous vehicles (AVs) but has broad robotic applications, e.g., elder-assistive robots [61] that fetch diverse items [23] should address LT3D. To study LT3D, we establish rigorous evaluation protocols that allow for partial credit to better diagnose 3D detectors. We propose several algorithmic innovations to improve LT3D, including a group-free detector head, a hierarchical loss that promotes feature sharing across long-tailed classes, and present a detailed exploration of multi-modal late-fusion for LT3D. We find that 2D RGB detectors are better suited for late-fusion, matching projected 3D LiDAR detections on the 2D image-plane outperforms matching 2D RGB detections inflated to 3D, and score calibration and probabilistic fusion notably improve performance. Our simple multi-modal late-fusion achieves state-of-the-art LT3D performance on the nuScenes benchmark, improving over prior works by 5.9 mAP over all long-tailed distributed classes.
Appendix A Implementation Details
We use the open-source codebase det3d mmcv, mmdetection and mmdetection3d to train FCOS3D, CenterPoint, and TransFusion. We use the first-party implementations for other detectors, following the training schedule proposed by each respective paper. By default, we train the 2D RGB detector with 2D bounding boxes derived from nuScenes’ 3D annotations and additionally train with 2D bounding boxes from nuImages where denoted. Our 2D RGB detectors, YOLOV7 and DINO, are pre-trained on the ImageNet [13] and COCO [36] datasets. We follow the default training recipes for both YOLOV7 and DINO respectively. We describe important implementation details below.
@@ Appendix A Implementation Details
  • Training Losses. We use the sigmoid focal loss (for recognition) [35] and L1 regression loss (for localization) below. Existing works also use the same losses but only with fine labels; we apply the loss to both coarse and fine labels. Concretely, our loss function for CenterPoint is as follows: $L=L_{HM}+\lambda L_{REG}$, where $L_{HM}=\sum_{i=0}^{C}SigmoidFocalLoss(X_{i},Y_{i})$ and $L_{REG}=|X_{BOX}-Y_{BOX}|$, where $X_{i}$ and $Y_{i}$ are the $i^{th}$ class’ predicted and ground-truth heat maps, while $X_{BOX}$ and $Y_{BOX}$ are the predicted and ground-truth box attributes. Without our hierarchical loss, CC=18. With our hierarchical loss, CC=22 (18 fine grained + 3 coarse + 1 object class). $\lambda$ is set to 0.25. Modifications for other detectors similarly follow.
  • Optimization. We train all LiDAR-only detectors for 20 epoch using an AdamW optimizer and a cyclic learning rate. We adopt a basic set of data augmentations, including global 3D tranformations, flip in BEV, and point shuffling during training. We train our model with 8 RTX 3090 GPUs and a batch size of 1 per GPU. The training noise (from random seed and system scheduling) is << 1% of the accuracy (standard deviation normalized by the mean).
@@
Appendix B State-of-the-Art Comparison on Argoverse 2
We present results on the large-scale Argoverse 2 (AV2) dataset developed for autonomous vehicle research. AV2 evaluates on 26 classes, which follow a long-tailed distribution. In Fig. 12, we plot the distribution on the left panel and show how we construct a semantic hierarchy on the right panel. Following prior work [46], we train on and evaluate detections up to 50m. As shown in Table 6, our main conclusions from nuScenes still hold for AV2. FCOS3D [76] yields poor performance on all classes, likely due to inaccurate depth estimates. CenterPoint performs considerably better, achieving high accuracy on classes with many examples. Notably, CenterPoint performs better on AV2’s rare classes (30.2 mAP) compared to nuScenes’s rare classes (3.5 mAP), likely because AV2 has more examples per-class in-the-tail. Lastly, our proposed late-fusion approach yields a 8.3 mAPimprovement over CenterPoint and a 3.9 mAP improvement over prior work. These new results on AV2 are consistent with those on nuScenes, demonstrating the general applicability of our approach.
Few       Medium       Many
Refer to caption Refer to caption
Method Modality All Many Medium Few
FCOS3D C 14.6 27.4 17.0 7.8
CenterPoint (CP) L 44.0 77.4 46.9 30.2
MMF(CP, FCOS3D) C + L 48.4 79.0 51.4 35.3
MMLF(CP, DINO) C + L 52.3 89.4 54.2 38.7
Table 6: Comparison with the Argoverse 2 state-of-the-art. We present results on AV2 evaluated at 50m. The RGB-only 3D detector FCOS3D [76] achieves poor performance, likely due to inaccurate depth estimates. In contrast, the LiDAR-only detector CenterPoint (CP) [82] achieves strong performance on all classes. Our multi-modal fusion approach significantly improves over CenterPoint, achieving an 8.3 mAP improvement averaged over all classes. Importantly, late-fusion with 2D detector DINO [88] performs the best by a notable gain. These results on AV2 are consistent with those on nuScenes, demonstrating the general applicability of our approach.
Appendix C More Visualizations
@@
Appendix D Analysis on Transformer 3D Detectors
In the main paper, we evaluate all late-fusion methods by filtering LiDAR detections from CenterPoint with different RGB detectors for fair comparison to [46]. However, our late-fusion approach generalizes to other 3D LiDAR detectors as well. To this end, we evaluate our best-performing late-fusion configuration from the main paper with LiDAR detections from BEVFusion-L [40] and CMT-L [80] in Table 7.
@@ Appendix D Analysis on Transformer 3D Detectors
Table 7: Transformer-based 3D LiDAR detector. We evaluate the late-fusion performance of BEVFusion-L and CMT-L (LiDAR-only) with DINO, and find that our late fusion strategy improves over BEVFusion-L by 8.2 mAP overall and CMT-L by 15.9 mAP. This suggests that our late-fusion approach generalizes across different 3D detector architectures.
Appendix E Inference Runtime
FCOS3D BEVFormer CenterPoint (CP) TransFusion DeepInteraction MMLF(CP, DINO)
@@ Appendix E Inference Runtime
Table 8: Comparison of inference time. Our multi-modal late-fusion approach runs faster than existing methods including end-to-end multimodal detectors such as DeepInteraction. Note that although our approach runs two uni-modal detectors, we run them in parallel so that the detection stage maintains the same efficiency as CenterPoint (CP). Moreover, our late-fusion strategy introduces negligible computation in inference. We copy 3D detection performance (in mAP) from the main paper for reference. Results demonstrate that our simple multi-modal late-fusion approach resoundingly outperforms sophisticated state-of-the-art methods (e.g., DeepInteraction and TransFusion) with significantly less inference time. Nevertheless, all methods require implementation optimization towards real-time processing for real-world application.
Table 8 compares the inference runtime of our method against prior works on a single GPU. Although our multi-modal late-fusion (MMLF) methods such as MMLF(CenterPoint, DINO) run two uni-modal detectors, we run them in parallel. Therefore, our final detector maintains the same runtime as a typical uni-modal detector and is often faster than more sophisticated multimodal detectors like BEVFormer [43], TransFusion [2] and DeepInteraction [79]. However, running two detectors in parallel will require more memory usage. Importantly, late-fusion methods including MMLF(CenterPoint, DINO) and [46] have negligible overhead when fusing uni-modal detections, hence their runtime is the same as CenterPoint, which serves as the speed bottleneck. This work uses research-level code and can be further optimized for deployment on autonomous vehicles.
Appendix F Analysis on Class Grouping
@@
Adult $P$ CW Debris Child Stroller PO EV PM All
@@
Table 9: Our proposed group-free detector head architecture consistently outperforms grouping-based approaches on both the standard and LT3D  benchmarks. We note that sub-optimal grouping strategies (such as those adopted for LT3D) may yield significantly diminished performance, whereas optimized grouping strategies (such as those adopted for the standard setup) have comparable performance to the group-free approach. Note, TC is traffic-cone, CV is construction vehicle, MC is motorcycle, $P$ is pushable-pullable, CW is construction-worker, and PO is police-officer. We highlight Medium and Few classes in blue.
@@
We use the class groups proposed by prior works [84, 82] for the standard benchmark and adapt this grouping for LT3D. Our proposed group-free detector head architecture consistently outperforms grouping-based approaches on both the standard and LT3D benchmarks, as shown in Table 9. We note that sub-optimal grouping strategies (such as those adopted for LT3D) may yield significantly diminished performance, whereas optimized grouping strategies (such as those adopted for the standard setup) have comparable performance to the group-free approach. The group-free approach simplifies architecture design, while also providing competitive performance.
@@
Appendix G Analysis of Hierarchical Training
@@
Table 10: Different variants achieve similar performance. We note that other methods do improve accuracy in the tail by sacrificing performance in the head, suggesting that hybrid approaches that apply different techniques for head-vs-tail classes may further improve accuracy. Unlike [42, 74] which requires a strict label hierarchy, our approach is not limited to a hierarchy.
@@
Classic methods train a hierarchical softmax (in contrast to our simple approach of sigmoid focal loss with both fine and coarse classes), where one multiplies the class probabilities of the hierarchical predictions during training and inference [74]. We implemented such an approach, but found the training did not converge. Interestingly, [74] shows such a hierarchical softmax loss has little impact on long-tailed object detection (in 2D images), which is one reason they have not been historically adopted. Instead, we found better results using the method from [42] (a winning 2D object detection system on the LVIS [19] benchmark) which multiples class probabilities of predictions (e.g. $P_{CAR}=P_{OBJ}*P_{CAR}$) at test-time, even when such predictions are not trained with a hierarchical softmax. We tested three variants and compared it to our approach (which recall, uses only fine-grained class probabilities at inference). Table 10 compares their performance for LT3D.
Class name w/ YOLOV7 w/ DINO
Personal Mobility 12.5 17.3
Emergency Vehicle 10.4 16.2
Police Officer 10.7 20.9
Stroller 22.1 32.3
Child 6.4 7.6
Debris 5.2 5.9
Construction Worker 40.4 54.6
Bicycle 63.1 68.2
Motorcycle 72.0 74.1
Construction Vehicle 31.7 33.6
Bus 75.3 75.9
Pushable Pullable 46.2 54.2
Trailer 40.8 43.3
Truck 58.5 60.1
Traffic Cone 79.9 80.5
Barrier 74.2 74.3
Adult 85.8 86.0
Car 86.9 86.1
Few 11.2 16.7
Medium 52.8 57.7
Many 77.1 77.4
All 45.7 49.5
Table 11: Using a stronger 2D detector improves 3D detection by MMLF. We compare using DINO [88] vs. YOLOv7 [70] as the 2D detector, applied along with the 3D LiDAR detector CenterPoint [82] in our MMLF. Results demonstrate that using DINO, a stronger detector than YOLOv7, yields significantly better 3D detection performance on all classes, particularly on medium and few classes. Refer to Fig. 1 for the distribution of these classes.
  1. (a)
    Ours (e.g., Fine-class score only)
  2. (b)
    Object score * Fine-class score ([42], e.g. $P_{CAR}=P_{OBJ}*P_{CAR}$)
  3. (c)
    Coarse score * Fine-class score (Variant-1 of [42], e.g. $P_{CAR}=P_{VEHICLE}*P_{CAR}$)
  4. (d)
    Object score * Coarse score * Fine-class score (Variant-2 of [42], e.g. $P_{CAR}=P_{OBJ}*P_{VEHICLE}*P_{CAR}$)
Appendix H Per-Class Breakdown Results
Class name BevFormer Deepinteraction CMT MMLF(CenterPoint, DINO)
[43] [79] [80] (Ours)
Personal Mobility 0.7 2.5 4.1 19.3
Emergency Vehicle 0.2 1.5 6.9 21.4
Police Officer 0.9 0.0 7.3 25.6
Stroller 1.1 30.9 0.9 37.7
Child 3.9 6.0 4.7 8.8
Debris 1.9 6.8 4.6 6.8
Construction Worker 21.5 30.7 34.4 55.9
Bicycle 41.9 63.3 66.7 70.0
Motorcycle 47.0 76.2 76.3 75.9
Construction Vehicle 16.9 35.3 36.9 35.3
Bus 43.0 75.9 76.4 76.1
pushable Pullable 29.1 30.8 34.1 58.1
Trailer 21.4 45.4 46.5 44.6
Truck 37.1 63.2 65.2 60.6
Traffic Cone 59.9 71.8 80.1 81.5
Barrier 53.2 75.3 77.9 74.6
Adult 50.3 85.9 87.7 86.2
Car 60.8 84.9 88.6 86.3
Few 1.4 7.9 4.8 20.0
Medium 31.6 51.1 53.0 59.5
Many 52.3 76.2 79.9 77.9
All 27.3 43.7 44.4 51.4
Table 12: Per-class performance on nuScenes. Our approach MMLF fuses detections from the LiDAR-based 3D detector CenterPoint and the RGB-based 2D detector DINO, achieving the highest per-class AP on 10 out of 18 classes. It is worth noting that, compared to CMT, our approach improves construction worker by 21.5 AP, stroller by 36.8 AP, and pushable-pullable by 24.0 AP.
We compare recent multi-modal methods w.r.t per-class AP in Table 12. All multi-modal methods perform similarly on common classes but considerably worse on Medium and Few classes, highlighting the need for further investigation of LT3D by the research community. Notably, our late-fusion approach achieves 20.0 higher AP on pushable-pullable and 6.0 higher AP on stroller than prior work. In general, our late-fusion approach yields considerable improvement on classes with medium and few examples. Despite significant improvements in rare class detection accuracy, our approach detects child with a low 8 mAP, which is still much better than compared methods. We posit that it is difficult to differentiate child from adult due to perspective geometry, since a small child close to the camera looks similar to a tall adult far away from the camera.
Appendix I Analysis of 2D Detection Performance